Write 2.85 as a percent?
An appliance dealer marks up refrigerators 20% ( based on cost). If the cost of one model is $500, what should it's selling price be?

Found 2 solutions by checkley71, stanbon:
Answer by checkley71(8403)

to change a number to % just move the decimal 2 places to the right.
2.85=285%
-----------------------------------------------------
500*1.20=$600 is the marked up or selling price.

Answer by stanbon(75887)

Write 2.85 as a percent?
Percent means per 100; 100 must be in the denominator.
2.85[100/100] = 285/100 = 285%
----------------------------------------
An appliance dealer marks up refrigerators 20% ( based on cost). If the cost of one model is $500, what should it's selling price be?
selling price = 500 + 0.20*500
selling price = 500 + 100
selling price = $600
Cheers,
San H.